OpenStudy (anonymous):

Find all critical points of the function. (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.) f(t) = 7t − t^2 + 7

OpenStudy (campbell_st):

find the derivative and solve for t \[df(t)/dt = 7 -2t\] solving by letting the derivative = 0 then t = 7/2 sub into the original equation to find f(t) point is (7/2 19 1/4)